OpenStudy (anonymous):

Please help with this 1st order DE: y'=3x^2-(y/x); y(1)=5

OpenStudy (anonymous):

Use the formula! The solution is uniquely determined because of the initial value: \[ y \left( x \right) =1/4\,{\frac {3\,{x}^{4}+17}{x}} \]

OpenStudy (anonymous):

y' + (y/x) = 3x^2 integrating factor = e ^(INT(1/x) dx = e ^lnx = x yx = INT3x^2*x dx yx = 3x^4 / 4 + C x=1 , y = 5 5 = 3/4 + c c = 17/4 therefore y = (1/4) (3x^4 + 17) / x